Calculate the wavelength associated with cricket ball of mass of 150g travelling with a velocity of 25ms-1.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
wavelength lambda = h/ mv = 6.255 * 10-34 / (0.15 * 25)
 
solve this
 
hope it helps
 
Regards
Arun